# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 2.3 Finite-Gain $\mathcal{L}_{p}$ Stability and Contraction of Hierarchically Combined Systems
If (2.30) is perturbed by ${d}({x},t)\in\mathcal{L}_{pe}$ (i.e., $\|(d)_{\tau}\|_{\mathcal{L}_{p}}<\infty$ for $\tau\in\mathbb{R}_{\geq 0}$ and $p\in[1,\infty]$, see Sec. 1.3) and Theorem 2.1 holds, then (2.30) is finite-gain $\mathcal{L}_{p}$ stable with $p\in[1,\infty]$ for an output y=h(x,d,t){y}={h}({x},{d},t) with $\int^{Y_{1}}_{Y_{0}}\|\delta{y}\|\leq\eta_{0}\int^{\xi_{1}}_{\xi_{0}}\|\delta{x}\|+\eta_{1}\|{d}\|$, $\exists\eta_{0},\eta_{1}\geq 0$, i.e., $\forall\tau\in\mathbb{R}_{\geq 0}$ Chung et al. [2013]
@@
In keeping with Theorem 5.1 of Khalil [2002], (2.36) also implies the following relation for $M=\Theta^{\top}\Theta$:
@@
Since we have $\|{\Theta}^{-1}\|\leq 1/\sqrt{\underline{m}}$ for $M=\Theta^{\top}\Theta$, (2.51) can be obtained by using both (2.52) and the known bound of $\|\delta{y}\|$, thereby yielding a finite $\mathcal{L}_{p}$ gain independently of $\tau$. ISS can be guaranteed by Lemma 4.6 of [Khalil, 2002, p. 176], which states that exponential stability of an unperturbed system results in ISS. $\qed$
Theorems 2.12.6 are also applicable to the hierarchically combined system of two contracting dynamics due to Theorem 2.7 Slotine [2003], Slotine and Lohmiller [2001].
@@
Consider the following hierarchically combined system of two contracting dynamics:
@@
where $F=F_{00}$ and $F=F_{11}$ both satisfy the contraction condition (2.5), and suppose that it is subject to perturbation ${d}_{0},{d}_{1}]^{\top}$. Then the path length integral $V_{\ell,i}(t)=\int^{\xi_{1}}_{\xi_{0}}\|\delta{z}_{i}\|$ with $i=0,1$ between the original and perturbed dynamics trajectories, $\xi_{0}$ and $\xi_{1}$, respectively, verifies Lohmiller and Slotine [1998]
@@
where $\alpha_{i}=\sup_{x,t}|\lambda_{\mathrm{max}}({F}_{ii})|,~i=0,1$. Hence, the error bounds of $V_{\ell,0}(t)$ and $V_{\ell,1}(t)$ can be obtained using Theorem 2.4 if $\|{F}_{10}\|$ is bounded. In particular, if $\|{\Theta}_{i}d_{i}\|\leq\sqrt{\overline{m}_{i}}\bar{d}_{i}$, $i=0,1$, the relations (2.54) and (2.55) yield $V_{\ell,0}(t)\leq V_{\ell,0}(0)e^{-\alpha_{0}t}+(\sqrt{\overline{m}_{0}}\bar{d}_{0}/{\alpha_{0}})(1-e^{-\alpha_{0}t})$ as in (2.33), and
@@
Also, similar to Theorem 2.6, we have $\|(V_{\ell,0})_{\tau}\|_{\mathcal{L}_{p}}\leq V_{\ell,0}(0)\zeta_{0}+{\|({\Theta}_{0}{d}_{0})_{\tau}\|_{\mathcal{L}_{p}}}/{\alpha_{0}}$, and thus a hierarchical connection for finite-gain $\mathcal{L}_{p}$ stability can be established as follows Chung et al. [2013]:
@@
Applying the comparison lemma of Lemma 2.1 to (2.54) and (2.55), which follow from the differenital dynamics (2.53) with the condition (2.5), we get (2.56) as in Theorem 2.4. Similarly, (2.57) follows from obtaining $\|(V_{\ell,0})_{\tau}\|_{\mathcal{L}_{p}}$ by (2.52) using (2.54), and then recursively obtaining $\|(V_{\ell,1})_{\tau}\|_{\mathcal{L}_{p}}$ using (2.55) as in Theorem 2.6 Chung et al. [2013]. $\qed$
@@
As demonstrated in Example 2.6, the Lagrangian virtual system (2.24) is contracting with respect to $q$, having $q=\dot{\mathtt{q}}$ and $q=\dot{\mathtt{q}}_{r}=\dot{\mathtt{q}}_{d}-\Lambda(\mathtt{q}-\mathtt{q}_{d})$ as its particular solutions. Let $q_{0}=q$ for such $q$. The virtual system of $q_{1}$ which has $q_{0}^{\top},q_{1}^{\top}]^{\top}=[\dot{\mathtt{q}}^{\top},\mathtt{q}^{\top}]^{\top}$ and $\dot{\mathtt{q}}_{r}^{\top},\mathtt{q}_{d}^{\top}]^{\top}$ as its particular solutions is given as
@@
resulting in $\delta\dot{q}_{1}=\mathrm{I}\delta q_{0}-\Lambda\delta q_{1}$. Since the virtual system of $q_{0}$ is contracting as in Example 2.6 and the virtual system $\delta\dot{q}_{1}=-\Lambda\delta q_{1}$ is contracting in $q_{1}$ due to $\Lambda\succ 0$, (2.56) of Theorem 2.7 implies that the whole system (2.22) for $q_{0}^{\top},q_{1}^{\top}]^{\top}$ is hierarchically contracting and robust against perturbation in the sense of Theorem 2.7 ($F_{10}=\mathrm{I}$ in this case). Also, see Chung and Slotine [2009], Chung et al. [2013] for the hierarchical multi-timescale separation of tracking and synchronization control for multiple Lagrangian systems.
2.4 Contraction Theory for Discrete-time Systems
@@
The discrete version of robust contraction in Theorem 2.4 is given in the following theorem.
@@
Let ${x}(k)=x_{k}$ and $q(\mu,k)=q_{k}$ for any $k\in\mathbb{N}$ for simplicity. If there exists a uniformly positive definite matrix ${M}_{k}({x}_{k},k)={{\Theta_{k}}}({x}_{k},k)^{\top}{{\Theta_{k}}}({x}_{k},k)\succ 0,~\forall x_{k},k$, where $\Theta_{k}$} defines a smooth coordinate transformation of $\delta x_{k}$, i.e., $\delta{z}_{k}={\Theta_{k}}(x_{k},k)\delta{x}_{k}$, s.t. either of the following equivalent conditions holds for $\exists\alpha\in(0,1)$, $\forall x_{k},k$:
@@
Theorem 2.8 can be used with Theorem 2.4 for stability analysis of hybrid nonlinear systems Slotine et al. [2004], Lohmiller and Slotine [2000], Pham [2008], or with Theorem 2.5 for stability analysis of discrete-time stochastic nonlinear systems Tsukamoto and Chung [2021, 2019], Pham [2008]. For example, it is shown in Tsukamoto and Chung [2021] that if the time interval in discretizing (2.1) as (2.59) is sufficiently small, contraction of discrete-time systems with stochastic perturbation reduces to that of continuous-time systems.
We finally remark that the steady-state upper bounds of (2.33) in Theorem 2.4, (2.47) in Theorem 2.5, and (2.63) in Theorem 2.8 are all functions of $\overline{m}/\underline{m}$. This property is to be used extensively in Sec. 4 for designing a convex optimization-based control and estimation synthesis algorithm via contraction theory.
3 Robust Nonlinear Control and Estimation via Contraction Theory
As shown in Theorem 2.4 for deterministic disturbance and in Theorem 2.5 for stochastic disturbance, contraction theory provides explicit bounds on the distance of any couple of perturbed system trajectories. This property is useful in designing robust and optimal feedback controllers for a nonlinear system such as $\mathcal{H}_{\infty}$ control Basar and Bernhard [1995], Van der Schaft [1992], Isidori and Astolfi [1992], Doyle et al. [1989], Khargonekar and Rotea [1991], Xie and de Souza Carlos [1992], Scherer et al. [1997], Chen and Zhang [2004], Hinrichsen and Pritchard [1998], Zhang and Chen [2006], Manchester and Slotine [2018], which attempts to minimize the system $\mathcal{L}_{2}$ gain for optimal disturbance attenuation.
Since most of such feedback control and estimation schemes are based on the assumption that we know a Lyapunov function candidate, this section delineates one approach to solve a nonlinear optimal feedback control problem via contraction theory Tsukamoto and Chung [2021, 2019], thereby proposing one explicit way to construct a Lyapunov function and contraction metric for general nonlinear systems for the sake of robustness. This approach is also utilizable for optimal state estimation problems as shall be seen in Sec. 4.
@@ 3 Robust Nonlinear Control and Estimation via Contraction Theory
We consider control-affine nonlinear systems (3.1)–(3.3) in Sec. 3, 4, and 68. This is primarily because the controller design techniques for control-affine nonlinear systems are less complicated than those for control non-affine systems (which often result in $u$ given implicitly by $u=k(x,u,t)$ Wang et al. [2020, 2019]), but still utilizable even for the latter, e.g., by treating $\dot{u}$ as another control input (see Example 3.1), or by solving the implicit equation $u=k(x,u,t)$ iteratively with a discrete-time controller (see Example 3.2 and Remark 3.3).
@@
By using $\dot{u}$ instead of $u$ in (3.1) and (3.2), a control non-affine system, $\dot{x}=f(x,u,t)$, can be rewritten as
@@
One drawback of the technique in Example 3.1 is that we have to control $\dot{u}$ instead of $u$, which could be difficult in practice. In this case, we can utilize the following control non-affine nonlinear system decomposed into control-affine and non-affine parts:
@@
where $u^{*}$ is a stabilizing controller for the control-affine system $\dot{x}=f_{a}(x,t)+B_{a}(x,t)u^{*}$. Since solving such an implicit equation in (3.6) in real-time could be unrealistic in practice, we will derive a learning-based approach to solve it iteratively for unknown $r(x,u,t)$, without deteriorating its stability performance (see Lemma 9.1 and Theorem 9.2 of Sec. 9).
3.1 Overview of Nonlinear Control and Estimation
@@
SDC formulation (Theorem 4.2Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Dani et al. [2015], Tsukamoto and Chung [2019] CCM formulation (Theorem 4.6Manchester and Slotine [2017], Singh et al. [2017]
@@
3.1.1 Systems with Known Lyapunov Functions
@@
As discussed in Sec. 1.2, there are several nonlinear systems equipped with a known contraction metric/Lyapunov function, such as Lagrangian systems [Slotine and Li, 1991, p. 392], whose inertia matrix $\mathcal{H}(\mathtt{q})$ defines its contraction metric (see Example 2.6), or the nonlinear SLAM problem Tan et al. [2017], Dani et al. [2015] with virtual synthetic measurements, which can be reduced to an LTV estimation problem Tan et al. [2017]. Once we have a contraction metric/Lyapunov function, stabilizing control and estimation laws can be easily derived by using, e.g., Sontag [1983, 1989], Khalil [2002]. Thus, those dealing primarily with such nonlinear systems should skip this section and proceed to Part II of this paper (Sec. 59) on learning-based and data-driven control using contraction theory. Note that these known contraction metrics are not necessarily optimal, and the techniques to be derived in Sec. 3 and Sec. 4 are for obtaining contraction metrics with an optimal disturbance attenuation property Tsukamoto and Chung [2021, 2019].
3.1.2 Linearization of Nonlinear Systems

### v8

@@ 2.3. Finite-Gain $\mathcal{L}_{p}$ Stability and Hierarchical Contraction
$\mathcal{L}_{p}$ Robust ContractionThm:Robust_contraction If (2.30) is perturbed by ${d}({x},t)\in\mathcal{L}_{pe}$ (i.e., $\|(d)_{\tau}\|_{\mathcal{L}_{p}}<\infty$ for $\tau\in\mathbb{R}_{\geq 0}$ and $p\in[1,\infty]$, see Sec. 1.3) and Theorem LABEL:THM:Thm:contraction holds, then (2.30) is finite-gain $\mathcal{L}_{p}$ stable with $p\in[1,\infty]$ for an output y=h(x,d,t){y}={h}({x},{d},t) with $\int^{Y_{1}}_{Y_{0}}\|\delta{y}\|\leq\eta_{0}\int^{\xi_{1}}_{\xi_{0}}\|\delta{x}\|+\eta_{1}\|{d}\|$, $\exists\eta_{0},\eta_{1}\geq 0$, i.e., $\forall\tau\in\mathbb{R}_{\geq 0}$ Ref:phasesync
@@
In keeping with Theorem 5.1 of Khalil:1173048, (2.36) also implies the following relation for $M=\Theta^{\top}\Theta$:
@@
Since we have $\|{\Theta}^{-1}\|\leq 1/\sqrt{\underline{m}}$ for $M=\Theta^{\top}\Theta$, (2.51) can be obtained by using both (2.52) and the known bound of $\|\delta{y}\|$, thereby yielding a finite $\mathcal{L}_{p}$ gain independently of $\tau$. ISS can be guaranteed by Lemma 4.6 of [Khalil:1173048, p. 176], which states that exponential stability of an unperturbed system results in ISS. $\qed$
Theorems LABEL:THM:Thm:contractionLABEL:THM:Thm:Robust_contraction are also applicable to the hierarchically combined system of two contracting dynamics due to Theorem LABEL:THM:Thm:Robust_contraction_hierc Ref:contraction2, Ref:contraction5.
@@
Robust Hierarchical ContractionThm:Robust_contraction_hierc Consider the following hierarchically combined system of two contracting dynamics:
@@
where $F=F_{00}$ and $F=F_{11}$ both satisfy the contraction condition (2.5), and suppose that it is subject to perturbation ${d}_{0},{d}_{1}]^{\top}$. Then the path length integral $V_{\ell,i}(t)=\int^{\xi_{1}}_{\xi_{0}}\|\delta{z}_{i}\|$ with $i=0,1$ between the original and perturbed dynamics trajectories, $\xi_{0}$ and $\xi_{1}$, respectively, verifies Ref:contraction1
@@
where $\alpha_{i}=\sup_{x,t}|\lambda_{\mathrm{max}}({F}_{ii})|,~i=0,1$. Hence, the error bounds of $V_{\ell,0}(t)$ and $V_{\ell,1}(t)$ can be obtained using Theorem LABEL:THM:Thm:Robust_contraction_original if $\|{F}_{10}\|$ is bounded. In particular, if $\|{\Theta}_{i}d_{i}\|\leq\sqrt{\overline{m}_{i}}\bar{d}_{i}$, $i=0,1$, the relations (2.54) and (2.55) yield $V_{\ell,0}(t)\leq V_{\ell,0}(0)e^{-\alpha_{0}t}+(\sqrt{\overline{m}_{0}}\bar{d}_{0}/{\alpha_{0}})(1-e^{-\alpha_{0}t})$ as in (2.33), and
@@
Also, similar to Theorem LABEL:THM:Thm:Robust_contraction, we have $\|(V_{\ell,0})_{\tau}\|_{\mathcal{L}_{p}}\leq V_{\ell,0}(0)\zeta_{0}+{\|({\Theta}_{0}{d}_{0})_{\tau}\|_{\mathcal{L}_{p}}}/{\alpha_{0}}$, and thus a hierarchical connection for finite-gain $\mathcal{L}_{p}$ stability can be established as follows Ref:phasesync:
@@
Applying the comparison lemma of Lemma 2.1 to (2.54) and (2.55), which follow from the differenital dynamics (2.53) with the condition (2.5), we get (2.56) as in Theorem LABEL:THM:Thm:Robust_contraction_original. Similarly, (2.57) follows from obtaining $\|(V_{\ell,0})_{\tau}\|_{\mathcal{L}_{p}}$ by (2.52) using (2.54), and then recursively obtaining $\|(V_{\ell,1})_{\tau}\|_{\mathcal{L}_{p}}$ using (2.55) as in Theorem LABEL:THM:Thm:Robust_contraction Ref:phasesync. $\qed$
@@
Contraction in Lagrangian Systemsex:hiera As demonstrated in Example LABEL:EX:ex:lag_metric, the Lagrangian virtual system (2.24) is contracting with respect to $q$, having $q=\dot{\mathtt{q}}$ and $q=\dot{\mathtt{q}}_{r}=\dot{\mathtt{q}}_{d}-\Lambda(\mathtt{q}-\mathtt{q}_{d})$ as its particular solutions. Let $q_{0}=q$ for such $q$. The virtual system of $q_{1}$ which has $q_{0}^{\top},q_{1}^{\top}]^{\top}=[\dot{\mathtt{q}}^{\top},\mathtt{q}^{\top}]^{\top}$ and $\dot{\mathtt{q}}_{r}^{\top},\mathtt{q}_{d}^{\top}]^{\top}$ as its particular solutions is given as
@@
resulting in $\delta\dot{q}_{1}=\mathrm{I}\delta q_{0}-\Lambda\delta q_{1}$. Since the virtual system of $q_{0}$ is contracting as in Example LABEL:EX:ex:lag_metric and the virtual system $\delta\dot{q}_{1}=-\Lambda\delta q_{1}$ is contracting in $q_{1}$ due to $\Lambda\succ 0$, (2.56) of Theorem LABEL:THM:Thm:Robust_contraction_hierc implies that the whole system (2.22) for $q_{0}^{\top},q_{1}^{\top}]^{\top}$ is hierarchically contracting and robust against perturbation in the sense of Theorem LABEL:THM:Thm:Robust_contraction_hierc ($F_{10}=\mathrm{I}$ in this case). Also, see Ref:ChungTRO, Ref:phasesync for the hierarchical multi-timescale separation of tracking and synchronization control for multiple Lagrangian systems.
2.4. Contraction Theory for Discrete-time Systems
@@
The discrete version of robust contraction in Theorem LABEL:THM:Thm:Robust_contraction_original is given in the following theorem.
@@
Discrete-Time ContractionThm:discrete_contraction Let ${x}(k)=x_{k}$ and $q(\mu,k)=q_{k}$ for any $k\in\mathbb{N}$ for simplicity. If there exists a uniformly positive definite matrix ${M}_{k}({x}_{k},k)={{\Theta_{k}}}({x}_{k},k)^{\top}{{\Theta_{k}}}({x}_{k},k)\succ 0,~\forall x_{k},k$, where $\Theta_{k}$} defines a smooth coordinate transformation of $\delta x_{k}$, i.e., $\delta{z}_{k}={\Theta_{k}}(x_{k},k)\delta{x}_{k}$, s.t. either of the following equivalent conditions holds for $\exists\alpha\in(0,1)$, $\forall x_{k},k$:
@@
Theorem LABEL:THM:Thm:discrete_contraction can be used with Theorem LABEL:THM:Thm:Robust_contraction_original for stability analysis of hybrid nonlinear systems Ref:contraction_robot, Ref:contraction4, 4795665, or with Theorem LABEL:THM:Thm:robuststochastic for stability analysis of discrete-time stochastic nonlinear systems mypaperTAC, mypaper, 4795665. For example, it is shown in mypaperTAC that if the time interval in discretizing (2.1) as (2.59) is sufficiently small, contraction of discrete-time systems with stochastic perturbation reduces to that of continuous-time systems.
We finally remark that the steady-state upper bounds of (2.33) in Theorem LABEL:THM:Thm:Robust_contraction_original, (2.47) in Theorem LABEL:THM:Thm:robuststochastic, and (2.63) in Theorem LABEL:THM:Thm:discrete_contraction are all functions of $\overline{m}/\underline{m}$. This property is to be used extensively in Sec. 4 for designing a convex optimization-based control and estimation synthesis algorithm via contraction theory.
3. Robust Nonlinear Control and Estimation
As shown in Theorem LABEL:THM:Thm:Robust_contraction_original for deterministic disturbance and in Theorem LABEL:THM:Thm:robuststochastic for stochastic disturbance, contraction theory provides explicit bounds on the distance of any couple of perturbed system trajectories. This property is useful in designing robust and optimal feedback controllers for a nonlinear system such as $\mathcal{H}_{\infty}$ control Basar1995, 256331, 159566, 29425, 85062, 151101, 599969, 1259458, doi:10.1137/S0363012996301336, doi:10.1137/S0363012903423727, rccm, which attempts to minimize the system $\mathcal{L}_{2}$ gain for optimal disturbance attenuation.
Since most of such feedback control and estimation schemes are based on the assumption that we know a Lyapunov function candidate, this section delineates one approach to solve a nonlinear optimal feedback control problem via contraction theory mypaperTAC, mypaper, thereby proposing one explicit way to construct a Lyapunov function and contraction metric for general nonlinear systems for the sake of robustness. This approach is also utilizable for optimal state estimation problems as shall be seen in Sec. 4.
@@ 3. Robust Nonlinear Control and Estimation
We consider control-affine nonlinear systems (3.1)–(3.3) in Sec. 3, 4, and 68. This is primarily because the controller design techniques for control-affine nonlinear systems are less complicated than those for control non-affine systems (which often result in $u$ given implicitly by $u=k(x,u,t)$ vccm, WANG201944), but still utilizable even for the latter, e.g., by treating $\dot{u}$ as another control input (see Example LABEL:EX:ex:nonaffine1), or by solving the implicit equation $u=k(x,u,t)$ iteratively with a discrete-time controller (see Example LABEL:EX:ex:nonaffine2 and Remark 3.3).
@@
Control Non-Affine Systems Iex:nonaffine1 By using $\dot{u}$ instead of $u$ in (3.1) and (3.2), a control non-affine system, $\dot{x}=f(x,u,t)$, can be rewritten as
@@
Control Non-Affine Systems IIex:nonaffine2 One drawback of the technique in Example LABEL:EX:ex:nonaffine1 is that we have to control $\dot{u}$ instead of $u$, which could be difficult in practice. In this case, we can utilize the following control non-affine nonlinear system decomposed into control-affine and non-affine parts:
@@
where $u^{*}$ is a stabilizing controller for the control-affine system $\dot{x}=f_{a}(x,t)+B_{a}(x,t)u^{*}$. Since solving such an implicit equation in (3.6) in real-time could be unrealistic in practice, we will derive a learning-based approach to solve it iteratively for unknown $r(x,u,t)$, without deteriorating its stability performance (see Lemma 9.1 and Theorem LABEL:THM:Thm:neurallander of Sec. 9).
3.1. Overview of Nonlinear Control and Estimation
@@
SDC formulation (Theorem LABEL:THM:Thm:CV-STEMmypaperTAC, ncm, nscm, Ref:Stochastic, mypaper CCM formulation (Theorem LABEL:THM:Thm:ccm_cvstemccm, 7989693
@@
3.1.1. Systems with Known Lyapunov Functions
@@
As discussed in Sec. 1.2, there are several nonlinear systems equipped with a known contraction metric/Lyapunov function, such as Lagrangian systems [Ref_Slotine, p. 392], whose inertia matrix $\mathcal{H}(\mathtt{q})$ defines its contraction metric (see Example LABEL:EX:ex:lag_metric), or the nonlinear SLAM problem doi:10.1177/0278364917710541, Ref:Stochastic with virtual synthetic measurements, which can be reduced to an LTV estimation problem doi:10.1177/0278364917710541. Once we have a contraction metric/Lyapunov function, stabilizing control and estimation laws can be easily derived by using, e.g., doi:10.1137/0321028, SONTAG1989117, Khalil:1173048. Thus, those dealing primarily with such nonlinear systems should skip this section and proceed to Part II of this paper (Sec. 59) on learning-based and data-driven control using contraction theory. Note that these known contraction metrics are not necessarily optimal, and the techniques to be derived in Sec. 3 and Sec. 4 are for obtaining contraction metrics with an optimal disturbance attenuation property mypaperTAC, mypaper.
3.1.2. Linearization of Nonlinear Systems